OpenStudy (anonymous):

find the range of values of k if the equation (2-3k)x2 + (4-k)x+2=0 has two different roots

OpenStudy (anonymous):

(2-3k)x^2 + (4-k)x+2=0 delta = (4-k)^2 - 8(2-3k) > 0 = k^2 - 8k + 16 - 16 + 24k >0 = k^2 + 16k >0 = k ( k + 16) >0 Thus k < 0 or k > 16

Directrix (directrix):

Given k ( k + 16) >0, what about the case where k< 0 AND (k + 16) <0?

OpenStudy (anonymous):

k<-16 or k>0